A diagonal of a square is 4 m. Find a side of the

square.



1
Expert's answer
2021-02-02T05:01:43-0500

We have that a diagonal of a square is "d=4\ \text{m}". Let "a" be the side of this square.

Using Pythagorean Theorem:

"a^2+a^2=d^2"

"2a^2=16"

"a^2=8"

"a=2\sqrt{2}"


Answer: the side of the square is "2\sqrt {2}\ \text{m}" .